# Decoding Neurodevelopmental Dynamics in Pediatric High-Grade Gliomas: Divergences in Transcriptomic Network Complexity from Healthy Neural Stem Cell Signatures

### v1

@@ 1 Introduction
Pediatric high-grade gliomas (pHGGs), such as diffuse midline gliomas (DMGs, or DIPGs in previous nomenclature) and glioblastomas (GBM) are complex adaptive ecosystems and are among the leading causes of cancer-related deaths in Western children[1, 2]. They exhibit a median survival of 1–2 years, primarily due to their invasive nature, intratumoral (molecular) heterogeneity, limited treatment penetration across the blood-brain barrier, high rates of recurrence, invasion, and therapy resistance, driven by multiscale regulatory networks [2, 3]. Much of the resilience and aggressiveness are attributed to their phenotypic plasticity, i.e., the adoption of a fluid spectrum of stem-like or progenitor-like phenotypic states with state-transitions driving disease progression [4, 5, 6, 3]. Plasticity acts as the ”evolvability engine” and cognitive scaffold shaping cell fate cybernetics, giving rise to emergent, maladaptive behaviors [7]. Despite debated differences in cells of origin, including developmentally arrested progenitor states, complex attractor dynamics from single-cell multiomics reveal a fluid, hybrid identity with a teleonomic bias toward specific terminal fates [8, 9, 10, 7]. However, these cell fate trajectories remain incomplete due to differentiation blockade, preventing attainment of their teleonomic endpoints (cell fate commitments) [8, 9, 10, 7].The behavioral patterns emerging from these arrested developmental processes, such as dysregulated cell fate decisions, are complex dynamical systems driven by plasticity networks [10].
@@
Achieving this requires complex dynamical systems approaches integrated with predictive causal inference algorithms and network science to forecast cell fate cybernetics (i.e., decision-making) [10] Complex systems theory, or complexity science, is the study of emergent (collective) behaviors in multiscale, interconnected dynamical systems, encompassing complex networks and nonlinear dynamics [12, 13]. Complexity science frames cancer ecosystems as ”processes” of multi-scale, system-level interactions and dynamic relationships, where the whole exhibits emergent behaviors beyond the sum of its parts[14, 7]. Through this framework, forecasting cell fate decisions and emergent behavioral patterns from plasticity networks in cancer ecosystems parallels weather forecasting in complex, chaotic systems [12, 13]. This perspective necessitates transdisciplinary frameworks from nonlinear dynamics, such as chaos theory, bifurcations, attractor dynamics, criticality, and network theory, to capture phase transitions and behavioral patterns in cancer networks [10]. As such, algorithmic information dynamics (AID) serves as a complexity-informed causal inference framework, to decode and predict the attractor dynamics underlying cancer state-transitions [15, 10]. AID provides a computational medicine toolkit for reconstructing transition and bifurcation events, mapping attractor dynamics, and inferring plasticity signatures [15].
In this study, we trace plasticity to epigenetic and developmental (lineage) plasticity by identifying plasticity signatures that steer and regulate cell fate transitions in glioma ecosystems. Rather than following the orchestrated developmental processes and differentiation hierarchies of healthy morphogenesis (pattern formation), pHGGs navigate a decentralized, multilineage attractor landscape that permits adaptive state-transitions and intratumoral heterogeneity, with a potential bias toward NPC/neuronal-like identities (see Supplementary Information).
@@ Discussion
Recent studies reveal that gliomas—particularly diffuse midline gliomas (DMGs) and DIPGs, as well as adult counterparts—integrate into neural circuits by forming neuron-to-glioma synapses, notably with GABAergic and glutamatergic neurons. These synaptic inputs promote malignant traits such as proliferation via depolarization-induced calcium signaling and neurotransmitter receptor signaling, representing hijacked neuroplasticity mechanisms [36, 37, 38]. Our findings suggest that beyond the emergence of invasive phenotypes, this synaptic integration reflects a deeper teleonomy (i.e., purposive, goal-directed behaviors) and semiotics (i.e., meaning-making processes) in glioma ecologies. They seem to suggest cell fate reprogramming toward neuron-like identities within the collective intelligence of glioma ecosystems throughout disease progression, as supported by transcriptomic enrichment of neuronal differentiation markers herein, suggesting that glioma fate cybernetics may be teleonomically biased toward neuronal functional states [23, 39]. These findings position glioma plasticity regulators within a (brain) neuron-glia-immune interface, as a therapeutic window to reprogram glioma ecosystems by targeting maladaptive neurotransmission dynamics.
@@
In contrast, GRIK3 and GRM8, are excitatory neuronal markers associated with neurodevelopmental disorders, and their emergence as differentially expressed or BDM signatures suggest a transient or hybrid neuronal identity, further highlighting the aberrant differentiation trajectories within glioma ecosystems. Furthermore, these transcriptional signatures support the emerging paradigm of systems medicine, where neuromodulators, and neurotransmitters along the neuroendocrine and brain-gut-immune axes play active cybernetic roles in steering and shaping glioma behaviors through synaptic feedback loops [43, 44, 45, 46, 37, 47]. Many studies have shown that brain tumors, in adults and children, such as glioblastomas exist within and exploit the complex neural circuitry of electrochemical and synaptic integration, where neuronal activity promotes tumor growth through both paracrine signaling and direct synaptic input [43, 44, 45, 46, 37, 47].For instance, recent studies show that tumor-promoting GABAergic neuron-to-glioma synapses in H3K27M-altered DMGs, where GABAA receptor-mediated depolarization drives glioma proliferation—unlike in hemispheric HGGs. Clinically used drugs like lorazepam were shown to worsen DMG growth, while levetiracetam suppressed it, revealing neuroanatomical location-specific and subtype-specific neurophysiological mechanism with therapeutic implications [48, 36]. This opens the door for precision neuropsychiatric therapies to be repurposed for treating gliomas, offering novel, systems targeted interventions based on their neurotransmissive and neural circuit-level dysregulation.

### v2

# Complexity-Informed Causal Modeling of Neurodevelopmental Trajectories in Pediatric High-Grade Gliomas: Divergences from Neural Stem Cell Signatures

@@ 1 Introduction
Pediatric high-grade gliomas (pHGGs), such as diffuse midline gliomas (DMGs, or DIPGs in previous nomenclature) and glioblastomas (GBM) are complex adaptive ecosystems and are among the leading causes of cancer-related deaths in Western children[1, 2]. They exhibit a median survival of 1–2 years, primarily due to their maladaptive behaviors, including, intratumoral (molecular) heterogeneity, limited treatment penetration across the blood-brain barrier, high rates of recurrence, invasion, and therapy resistance, driven by multiscale regulatory networks [2, 3]. Much of the resilience and aggressiveness are attributed to their phenotypic plasticity, i.e., the adoption of a fluid spectrum of stem-like or progenitor-like phenotypic states with state-transitions driving disease progression [4, 5, 6, 3]. Plasticity acts as the ”evolvability engine” and cognitive scaffold shaping cell fate cybernetics, giving rise to emergent, maladaptive behaviors [7]. Despite debated differences in cells of origin, including developmentally arrested progenitor states, complex attractor dynamics from single-cell multiomics reveal a fluid, hybrid identity with a teleonomic bias toward specific terminal fates [8, 9, 10, 7]. However, these cell fate trajectories remain incomplete due to differentiation blockade, preventing attainment of their teleonomic endpoints (cell fate commitments) [8, 9, 10, 7].The behavioral patterns emerging from these arrested developmental processes, such as dysregulated cell fate decisions, are complex dynamical systems driven by plasticity networks [10].
@@
Achieving this requires complex dynamical systems approaches integrated with predictive causal inference algorithms and network science to forecast cell fate cybernetics (i.e., decision making) [10]. Complex systems theory, or simply complexity science, is the study of emergent collective behaviors in interconnected dynamical systems, encompassing complex networks and nonlinear dynamics [12, 13]. It frames cancer ecosystems as network processes of multi-scale, system-level interactions and dynamic relationships, where the whole exhibits adaptive behaviors beyond the sum of its parts [14, 7]. In such systems, every agent or sub-part encodes traces of the whole, and the whole is interdependent and interwoven; identity emerges only in relation to other parts, through reciprocal agent–environment interactions as feedback loops. Thence, cancer ecology can only be understood as units of interactions, irreducible to isolated components or agents.
Through this framework, forecasting cell fate decisions and emergent behavioral patterns from plasticity networks parallels weather forecasting in complex, chaotic systems [12, 13]. This perspective necessitates transdisciplinary frameworks from nonlinear dynamics, such as chaos theory, bifurcations, attractor dynamics, criticality, and network theory, to capture phase transitions and behavioral patterns in cancer networks [10]. As such, algorithmic information dynamics (AID) serves as a complexity-informed causal inference framework, to decode and predict the attractor landscapes underlying cancer state-transitions [15, 10]. AID provides a computational medicine toolkit for reconstructing transition and bifurcation events, mapping attractor dynamics, and inferring plasticity signatures [15]. By network perturbation analysis, algorithmic or graph-complexity methods decode the causal information dynamics of transcriptional processes driving cell-fate trajectories[15].
In this study, we trace glioma plasticity to epigenetic and developmental (lineage) plasticity by identifying plasticity signatures that steer and regulate cell fate transitions in glioma ecosystems. Rather than following the orchestrated developmental processes and differentiation hierarchies of healthy morphogenesis (pattern formation), pHGGs navigate a decentralized, multilineage attractor landscape that permits adaptive state-transitions and intratumoral heterogeneity, with a potential bias toward NPC/neuronal-like identities (see Supplementary Information).
@@ Discussion
Recent studies reveal that gliomas—particularly diffuse midline gliomas (DMGs) and DIPGs, as well as adult counterparts—integrate into neural circuits by forming neuron-to-glioma synapses, notably with GABAergic and glutamatergic neurons. These synaptic inputs promote malignant traits such as proliferation via depolarization-induced calcium signaling and neurotransmitter receptor signaling, representing hijacked neuroplasticity mechanisms [36, 37, 38]. Our findings suggest that beyond the emergence of invasive phenotypes, this synaptic integration reflects a deeper (intrinsic) teleonomy (i.e., purposive, goal-directed behaviors) and semiotics (i.e., meaning-making processes) in glioma ecologies. They seem to suggest cell fate reprogramming toward neuron-like identities within the collective intelligence of glioma ecosystems throughout disease progression, as supported by transcriptomic enrichment of neuronal differentiation markers herein, suggesting that glioma fate cybernetics may be teleonomically biased toward neuronal functional states [23, 39]. These findings position glioma plasticity regulators within a (brain) neuron-glia-immune interface, as a therapeutic window to reprogram glioma ecosystems by targeting maladaptive neurotransmission dynamics.
@@
In contrast, GRIK3 and GRM8, are excitatory neuronal markers associated with neurodevelopmental disorders, and their emergence as differentially expressed or BDM signatures suggest a transient or hybrid neuronal identity, further highlighting the aberrant differentiation trajectories within glioma ecosystems. Furthermore, these transcriptional signatures support the emerging paradigm of systems medicine, where neuromodulators, and neurotransmitters along the neuroendocrine and brain-gut-immune axes play active cybernetic roles in steering and shaping glioma behaviors through synaptic feedback loops [43, 44, 45, 46, 37, 47]. Many studies have shown that brain tumors, in adults and children, such as glioblastomas exist within and exploit the complex neural circuitry of electrochemical and synaptic integration, where neuronal activity promotes tumor growth through both paracrine signaling and direct synaptic input [43, 44, 45, 46, 37, 47].For instance, recent studies show that tumor-promoting GABAergic neuron-to-glioma synapses in H3K27M-altered DMGs, where GABAA receptor-mediated depolarization drives glioma proliferation—unlike in hemispheric HGGs. Clinically used drugs like lorazepam were shown to worsen DMG growth, while levetiracetam suppressed it, revealing neuroanatomical location-specific and subtype-specific neurophysiological mechanism with therapeutic implications [48, 36]. This opens the door for precision neuropsychiatric therapies to be repurposed for treating gliomas, offering novel, systems targeted interventions based on their neurotransmissive and neural circuit-level dysregulation. Importantly, while previous studies emphasize glioma–neuron synaptic connectomes, our presented findings reveal neurotransmitter and synaptic gene expression signatures intrinsic to gliomas themselves, embedded within their single-cell fate decisions and transcriptome-level circuitry.